# Rank of elliptic curves and class groups of real quadratic fields

### v1

@@ 1. Introduction
where $a,b$ are rational integers and $4a^{3}+27b^{2}\neq 0$. Let the Mordell-Weil group $E({\mathbb{Q}})$ has positive rank and there exists a rational point on $E$, except the rank point. Then for an infinite number of prime numbers $p$, $p^{3}+ap+b$ is a square free negative integer and the class group of ${\mathbb{Q}}(\sqrt{p^{3}+ap+b})$ contains a non-trivial element which is obtained by specializing a generator of the rank part of $E({\mathbb{Q}})$.
Here, the existence of infinitely many primes such that $p^{3}+ap+b$ is square exists by the Siegel’s theorem that the number of integral points on an elliptic curve with integer coefficients is finite, so the number of solutions of the equation $y^{2}=p^{3}+ap+b$ is finite and hence there are infinitely many primes $p$ such that $p^{3}+ap+b$ is square-free.
@@
For the proof of this theorem we consider the technique of Chow schemes or Hilbert schemes on a scheme or variety defined over ${\rm{Spec}}({\mathbb{Z}})$. Let $E_{{\mathbb{Z}}}$ be a smooth integral model of the curve $E$ defined over ${\rm{Spec}}({\mathbb{Z}})$. Taking a point on the elliptic curve as a Weil divisor on the curve, that is, using an isomorphism $E\cong{\rm{Pic}}^{0}(E)$ we have such a realization, we spread the Weil divisor on $E_{{\mathbb{Z}}}$ to obtain a divisor on the fixed integral model of $E$ and then specialize this divisor at a prime integer p-p of $Spec({\mathbb{Z}})$ to obtain an element of the ideal class group of a number field given by ${\mathbb{Q}}(\sqrt{p^{3}+ap+b})$. The only thing is that if we start with an element of infinite order of the elliptic curve defined over $\mathbb{Q}$} we have to prove that the ideal class element that it produces in ${\mathbb{Q}}(\sqrt{p^{3}+ap+b})$ is non-principal.

### v2

@@ 1. Introduction
where $a,b$ are rational integers and $4a^{3}+27b^{2}\neq 0$. Let the Mordell-Weil group $E({\mathbb{Q}})$ has positive rank and there exists a rational point on $E$, except the rank point. Then for an infinite number of prime numbers $p$, $p^{3}+ap+b$ is a square free positive integer and the class group of ${\mathbb{Q}}(\sqrt{p^{3}+ap+b})$ contains a non-trivial element which is obtained by specializing a generator of the rank part of $E({\mathbb{Q}})$.
Here, the existence of infinitely many primes such that $p^{3}+ap+b$ is a non-square exists by the Siegel’s theorem that the number of integral points on an elliptic curve with integer coefficients is finite, so the number of solutions of the equation $y^{2}=p^{3}+ap+b$ is finite and hence there are infinitely many primes $p$ such that $p^{3}+ap+b$ is non-square.
@@
For the proof of this theorem we consider the technique of Chow schemes or Hilbert schemes on a scheme or variety defined over ${\rm{Spec}}({\mathbb{Z}})$. Let $E_{{\mathbb{Z}}}$ be a smooth integral model of the curve $E$ defined over ${\rm{Spec}}({\mathbb{Z}})$. Taking a point on the elliptic curve as a Weil divisor on the curve, that is, using an isomorphism $E\cong{\rm{Pic}}^{0}(E)$ we have such a realization, we spread the Weil divisor on $E_{{\mathbb{Z}}}$ to obtain a divisor on the fixed integral model of $E$ and then specialize this divisor at a prime integer $p$ of $Spec({\mathbb{Z}})$ to obtain an element of the ideal class group of a number field given by ${\mathbb{Q}}(\sqrt{p^{3}+ap+b})$. The only thing is that if we start with an element of infinite order of the elliptic curve defined over $\mathbb{Q}$} we have to prove that the ideal class element that it produces in ${\mathbb{Q}}(\sqrt{p^{3}+ap+b})$ is non-principal.